# Energy-Conscious LLM Decoding: Impact of Text Generation Strategies on GPU Energy Consumption

### v1

@@ 4.1.1 Key Findings:
Across all tasks, AD consistently delivers outstanding energy efficiency while demonstrating near-competitive text generation quality. In contrast, widely adopted beam-based deterministic methods offer strong generative quality but also consume disproportionately high amounts of energy, calling into question their default and unnecessary use for every application. Top‐p, Top‐k, and temperature‐based sampling methods exhibit similar power consumption across tasks, resulting in very similar efficiency ratios. Within an acceptable margin, all stochastic methods—when tuned to their optimal hyperparameters—consume essentially the same amount of energy across all tasks. This consistent pattern suggests that, despite their different approaches to sampling, they impose comparable computational overhead when generating output, making them viable alternatives for each other in scenarios where a stable trade‐off between energy efficiency and performance is desired across the selected tasks. Contrastive Search stands out in certain tasks, yet its significant energy draw negatively affect its overall energy efficiency. Except for GSM8k, DoLa consistently ranks among the worst in terms of efficiency ratio. Although this newly introduced method aims to reduce hallucinations and improve truthfulness in LLM outputs, it does not offer high generation quality in our chosen tasks and remains one of the highest energy‐consuming approaches. Finally, the overall energy consumption pattern in each decoding strategy seems to remain consistent across benchmarks. This creates opportunity to optimize decoding approaches in light of both quality requirements and energy considerations.
@@ 4.2 Hyperparameter Sensitivity Analysis
We also employed the Sharpe Ratio [43], a classic statistical metric from finance, to incorporate the recorded energy variations observed in each run for a given hyperparameter within our analysis. Mathematically expressed as $\frac{\mu-r_{f}}{\sigma}$, Sharpe Ratio measeures how well the average return of an investment compensates the volitality (risks) of returns. This reward-to-variability metric can be adapted in our analysis by treating the average of text generation qualities as the mean return of investments and the variability (standard deviation) of energy consumptions across all hyperparameter runs as the risks for each decoding strategy. This approach provides us a single metric to identify decoding strategies that are less sensitive to hyperparameter changes while simultaneously rewarding strong generative quality and penalizing energy fluctuations. A higher SR ratio signifies that the in all of our experiment runs for a specific decoding method, it consistently delivers good quality while minimizing energy consumption variations. Since we are ranking all decoding strategies within each task, we consider the risk‐free rate to be zero. Table 2 presents the RSD of text generation quality and average energy consumption across all hyperparameter values for each decoding strategy, along with the SR between quality mean and energy consumption variations across all hyperparameter runs.
Interestingly, beam-based methods exhibit relatively low RSD in BLEU scores, indicating stable generation quality across hyperparameter changes in both translation tasks. However, in both tasks, we observe a similar pattern where CS and DoLa methods show the highest variations in quality and energy, highlighting their strong sensitivity to hyperparameter tuning. Stochastic methods like Typical and Min‐p, and Top-p sampling generally remain more stable in both quality and energy, resulting in competitive Sharpe Ratios. In contrast, we find that Temperature sampling has significantly higher RSD values for both quality and energy compared to other stochastic techniques.
In GSM8K, we observe a similar pattern with CS and Temperature sampling methods exhibiting the highest variations in both quality and energy across their hyperparameters. Interestingly, AD shows one of the highest RSD values for quality, while its energy consumption remains relatively stable. In contrast, beam-based methods display the opposite trend, with minimal variations in quality but the highest fluctuations in energy across their hyperparameters. Another notable observation is that DoLa maintains extremely low energy variation, resulting in the highest SR ratio among all methods. Similar to translation tasks, most stochastic methods achieve a good balance between performance and energy stability, leading to high SR ratios.
Across both Python and JavaScript code summarization tasks, we observe a similar pattern where CS and Temperature sampling exhibit the highest variation in generation quality across hyperparameters. In contrast, DoLa remains significantly more stable, leading to the highest SR ratio in Python summarization. Methods like Beam and DBS experience considerable energy fluctuations while maintaining moderate variability in generation quality. Similar to previous tasks, many stochastic approaches demonstrate relatively low RSD in energy consumption and moderate RSD in quality, resulting in high SR in code summarization tasks.
4.2.1 Key Findings
Across all tasks, we observe that, except for Temperature sampling, the other stochastic decoding methods achieve some of the highest SR ratios, maintaining good quality with minimal energy fluctuations across different hyperparameter choices. This suggests that these methods are less sensitive to hyperparameter tuning and are more energy-conscious relative to the quality they produce. The reason Temperature sampling exhibits significant variation in both energy consumption and quality is the inclusion of high-temperature values in our choices. Higher temperatures $\tau>1$ increase randomness in token selection, reducing output fluency and coherence. This quality loss is further illustrated in Figure 2. Across all selected tasks, the model struggled to enhance diversity without sacrificing accuracy, leading to a noticeable drop in quality and increased energy consumption, particularly in translation and math problem-solving tasks. This is particularly interesting because, while Temperature sampling was among the methods with the highest efficiency ratio in Table 1, the RSD results in Table 2 reveals that it has the greatest variation in both quality and energy across different hyperparameters. This underscores the importance of hyperparameter optimization when using this method.
@@ 4.2.1 Key Findings
In some cases, such as beam-based methods in the math problem-solving task, we observe significant energy fluctuations despite minimal variation in quality. Conversely, in code summarization tasks, Temperature sampling exhibits the opposite trend, with substantial quality variation but stable energy usage. This proves that variations in quality and energy across hyperparameters do not always correlate.
@@
Another notable finding is that, while in Section 4.1 we identified AD as the most energy-efficient method in its best hyperparameter configuration for each task, We observe that only in the German-English translation task does it rank among the best methods with the highest SR ratio. In other cases, it exhibits variations in both generation quality and energy consumption across its selected hyperparameters, implying that AD performs optimally only when properly tuned in some text generation tasks. Additionally, for beam-based methods, we observe considerable energy fluctuations when increasing the beam width in code summarization and math problem-solving tasks (Figure 3). However, in translation tasks, energy consumption remains relatively stable, and the quality does not vary significantly. Both of these findings highlight the importance of task-specific hyperparameter tuning for certain decoding strategies.
@@
Contrastive Search stands out as the most sensitive to hyperparameter changes among all methods, with $k=25$ exhibiting significantly greater energy fluctuations than $k$ across different penalty $\alpha$ values. Larger $k$ values can lead to substantial shifts in both performance and energy consumption. Additionally, the impact of the $\alpha$ parameter varies considerably depending on the task, contributing to its high RSD values in both quality (up to 104.95%) and energy consumption (up to 53.45%).
@@
Lastly, we observe a unique pattern in the DoLA method. While Table 1 shows that it is among the methods with the highest energy consumption in its best hyperparameter setting, it also achieves some of the highest SR ratios in math problem-solving and translation tasks in Table 2. This indicates that, although DoLA may consume significant energy, its energy usage remains stable across different hyperparameters while still delivering acceptable (though not outstanding) generation quality in certain NLP tasks.
@@ 5 Limitations & Future Work
In this study, we confined our experiments to a relatively small-scale LLM, Qwen2.5 7B-Instruct, leaving the investigation of larger models and architectures for future work. While we carefully selected hyperparameters for our experiments, our approach was not exhaustive, leaving a broader range of parameter settings yet to be explored. Additionally, because LM-Evaluation-Harness framework did not offer benchmarking open-ended text generation tasks at the time of conducting this research, we focused on a subset of directed tasks that best suited our objectives. Recent decoding strategies research efforts significantly revolves around open-ended language generation, such as story creation and text continuation; thus, exploring how various decoding techniques influence energy consumption in such scenarios would be a valuable direction to pursue. We also did not combine different decoding strategies, suggesting another avenue for future research. Finally, our analysis focused on GPU energy consumption; however, future work could extend this scope to encompass other factors such as CPU power usage, memory and temperature, providing a more comprehensive understanding of how decoding strategies influence overall resource consumption.
@@ 6 Ethical Considerations and Societal Impact
7 Conclusion
In this study, we investigated the impact of decoding strategies and hyperparameter choices on both generative quality and GPU energy consumption during LLM inference across various text generation tasks. Through a comprehensive evaluation, we found that, in all tasks, the choice of decoding strategy significantly affects GPU energy consumption during inference, even in cases where it has minimal impact on output quality. Additionally, we observed that different decoding methods present distinct trade-offs between text quality and energy efficiency. While deterministic methods generally achieve top-tier quality in most generation tasks, they also tend to consume more energy and exhibit higher variability across their hyperparameters. In contrast, stochastic approaches appear to be more energy-efficient, often providing a better quality-to-energy ratio and demonstrating lower sensitivity to hyperparameter changes relative to their output quality. Overall, no single decoding strategy outperforms all others across every metric. Instead, the optimal choice depends on the specific task and whether the priority is maximizing generation quality, minimizing energy consumption, or achieving a stable balance between both. Our findings highlight the importance of carefully selecting decoding strategies, taking into account not only output quality but also computational and energy costs.
@@ Appendix A Appendix
A.1 Detailed Results
A.1.1 energy across hyperparameters:
In this section, we present figures illustrating how average energy consumption changes with varying hyperparameters for each decoding strategy. Additionally, The marker sizes in each graph represent generation quality, with larger markers indicating higher quality and smaller markers indicating lower quality.
Figure 2: The average energy consumption across different temperature values in Temperature sampling
Figure 3: The average energy consumption across different beam widths in Beam Search
Figure 4: The average energy consumption across different probability p in Top-p Sampling
Figure 5: The average energy consumption across different k values in Top-k Sampling
Figure 6: The average energy consumption across different DoLa layers. The numbers above each bar represent the corresponding generation quality
Figure 7: The average energy consumption across different typicality thresholds in Typical Sampling
Figure 8: The average energy consumption across different $\epsilon$ thresholds in Epsilon Sampling
Figure 9: The average energy consumption across different base probabilities in Min-p Sampling
Figure 10: The average energy consumption across different $\alpha$ values in Contrastive Search (k=5)
Figure 11: The average energy consumption across different $\alpha$ values in Contrastive Search (k=25)
Figure 12: The average energy consumption across different number of prompt tokens in Assisted Decoding
Table 3: Strategies and Hyperparameters. Each parameter was run 5 times to account for inference variations
strategies hyperparameter value
temperature {0.1, 0.3, 0.5, 0.7, 0.9, 1.0, 1,2, 1,5, 2}
epsilon {0.0003, 0.0006, 0.0009, 0.002, 0.004}
contrastive $\alpha:${0.2, 0.4, 0.6, 0.8, 1.0} $k:${5, 25}}
diverse beam {(4,2), (5,5), (10,2), (10,5)}
top_p {0.3, 0.5, 0.7, 0.8, 0.9, 0.95}
top_k {1, 5, 25, 50, 100, 250, 500}
min_p {0.02, 0.05, 0.1, 0.2, 0.3}
typical {0.2, 0.5, 0.8, 0.9, 0.95}
dola {low, high}
assisted {2, 5, 10}
beam search {2, 5, 10}
Table 4: Optimal Hyperparameters for the Decoding Strategies used in the experiments. The Greedy decoding method does not require any hyperparameters, so we represented it with a dash (-) in the table.

### v2

@@ 4.1.1 Overall Analysis of Decoding Strategies
Based on the average ranking of decoding strategies across all tasks and models, AD consistently has the lowest energy consumption, followed by Greedy, Min-p, Top-p, and Temperature sampling. Beam Search, CS, DBS, and DoLa rank as the worst methods in terms of energy consumption. For the average ranking of generative quality, Beam Search, DBS, Top-p, and Min-p are the best-performing methods, while Epsilon, DoLA, and Top-k are the worst-performing ones. Although beam-based deterministic methods offer strong generative quality, they consume disproportionately high amounts of energy, calling into question their unnecessary use. Within a reasonable margin, stochastic decoding methods, when tuned to their optimal hyperparameters, show similar power consumption across tasks. This consistent pattern suggests that, despite using different sampling approaches, they impose comparable computational overhead during output generation, making them viable alternatives for each other in cases where only energy efficiency matters. Contrastive Search performs well in some tasks, but its high energy usage lowers its overall efficiency ratio. DoLa consistently ranks among the worst in terms of efficiency ratio. Although this method aims to reduce hallucinations and improve truthfulness in LLM outputs, it does not provide strong generative quality in our tasks and remains one of the most energy-consuming techniques.
4.1.2 Per-Model Analysis of Decoding Strategies
Comparing the results of Qwen2.5-7B-Instruct with Llama3.1-8B-Instruct (Table 7 in Appendix), we see that in both translation tasks, the energy consumed by the two models across most decoding strategies is relatively close. However, since Llama3.1 generally achieves higher generative quality, its efficiency ratios are larger in most cases. Interestingly, in GSM8K, Qwen2.5-7B shows much better generation accuracy in many cases, but still has lower efficiency ratios because it consumes more energy on this task compared to Llama3.1. If accuracy is prioritized, Qwen2.5-7B is the better choice, but if energy consumption is the focus, Llama3.1 is superior. In coding tasks, Qwen2.5-7B performs better in both quality and energy use for most decoding methods. For Wikitext, we see a similar trend: Qwen2.5-7B achieves slightly higher MAUVE scores while also using less energy, resulting in better efficiency ratios.
Comparing the results of Qwen2.5-7B-Instruct with Qwen2.5-14B-Instruct (Table 8 in Appendix), we observe that although the 14B version achieves better generative performance in most tasks, its much higher energy consumption across all decoding methods leads to lower efficiency ratios compared to the 7B variant. On the Wikitext dataset, however, the 7B model delivers much higher MAUVE scores despite the 14B model consuming considerably more energy. This highlights that using larger models does not always guarantee better generative quality across all text generation scenarios.
Furthermore, looking at Figure 1 and comparing the RSD of GPU energy consumption across different decoding strategies in the three tested models, we observe that the deviation in energy can be large in specific tasks like GSM8K and HumanEval for Llama3.1-8B and Qwen2.5-14B compared to Qwen2.5-7B. This shows that models from different architectures or sizes can behave quite differently in certain text generation tasks.
Regarding the ranking of decoding strategies based on quality and energy consumption in each model, Beam Search is consistently the best-performing strategy for generative quality across all three tested models. Top-p and DBS also rank among the best after Beam Search, while DoLA, epsilon, and Top-k sampling are always the weakest methods. For energy consumption, AD consistently achieves the best ranking, with Greedy and Min-p also performing well in terms of energy efficiency. In contrast, CS, DBS, and DoLA rank the worst for energy usage.
4.1.3 Per-Task Analysis of Decoding Strategies
When evaluating the ranking of decoding strategies in each task, Beam Search and DBS consistently achieve the highest quality in both translation tasks and the HumanEval benchmark. In math problem-solving, Top-p and Beam Search perform best. DoLA, Epsilon, and Top-k rank the lowest in these tasks. In Wikitext, however, Top-k and Temperature sampling achieve the highest average ranks, while AD and DBS show the lowest rankings.
Regarding energy usage, AD consistently has the highest ranking across all tasks, while Greedy, Top-p, and Min-p are also among the best-performing strategies. In contrast, DBS, CS, DoLA, and Beam Search rank the lowest in terms of energy consumption.
Lastly, we observe that in specific tasks, such as German-to-English translation in Qwen2.5-14B or Llama3.1-8B, stochastic sampling strategies can sometimes consume more energy than Beam Search. This shows that energy usage is highly task- and model-dependent, and even sampling-based methods may use more energy than beam-based methods in certain cases.
@@ 4.2 Hyperparameter Sensitivity Analysis
We also employed the Sharpe Ratio [57], a classic statistical metric from finance, to incorporate the recorded energy variations observed in each run for a given hyperparameter within our analysis. Mathematically expressed as $\frac{\mu-r_{f}}{\sigma}$, Sharpe Ratio measeures how well the average return of an investment compensates the volitality (risks) of returns. This reward-to-variability metric can be adapted in our analysis by treating the average of text generation qualities as the mean return of investments and the variability (standard deviation) of energy consumptions across all hyperparameter runs as the risks for each decoding strategy. This approach provides us a single metric to identify decoding strategies that are less sensitive to hyperparameter changes while simultaneously rewarding strong generative quality and penalizing energy fluctuations. A higher SR ratio signifies that the in all of our experiment runs for a specific decoding method, it consistently delivers good quality while minimizing energy consumption variations. Since we are ranking all decoding strategies within each task, we consider the risk‐free rate to be zero. Table 3 presents the RSD of text generation quality and average energy consumption across all hyperparameter values for each decoding strategy, along with the SR between quality mean and energy consumption variations across all hyperparameter runs in Qwen2.5-7B-Instruct. The results for Llama3.1-8B-Instruct and Qwen2.5-14B-Instruct models are listed in Tables 9 & 10 in Appendix respectively.
Interestingly, Beam Search and DBS show relatively low RSD in quality for most tasks, especially in Translation and HumanEval across all the three tested models. However, they have relatively high RSD in energy, indicating stable generative performance but large variation in energy consumption when hyperparameters change. Although DoLA shows large variations in both quality and energy across its hyperparameters in most text generation tasks for the Qwen2.5 models, we observe that in the Llama3.1 model it remains impressively stable, with negligible change in either quality or energy. This highlights the point that the effect of hyperparameters on decoding strategies can differ greatly depending on the model family. Contrastive Search shows the highest variation in generation quality and energy usage across hyperparameters in all three models, especially when k is set to a large value, highlighting its strong sensitivity to hyperparameter tuning. Stochastic methods, except for Temperature sampling, are consistently among the strategies with the least variation in energy across different tasks and models. However, they show moderate fluctuations in generative quality across their hyperparameters. For Temperature sampling, we observe significantly high RSD values for both quality and energy in the Llama3.1 model. The reason Temperature scaling exhibits significant variation in both energy consumption and quality is the inclusion of high-temperature values in our hyperparameter choices. Higher temperatures $\tau>1$ increase randomness in token selection, reducing output fluency and coherence. Thus, the model struggles to enhance diversity without sacrificing accuracy, leading to a noticeable drop in quality and increased energy consumption. Surprisingly, in the Qwen2.5 variants, only quality shows large variation with temperature scaling, while energy remains relatively stable across the hyperparameter range. This highlights that large variations in generative quality do not always lead to significant fluctuations in energy consumption. Although in previous section we observed AD as the strategy with the lowest energy consumption in all tested scenarios, we find that it can show large variations in both quality and energy for tasks like GSM8K and HumanEval. In the other tasks we tested, both quality and energy remain stable across different hyperparameters for this deterministic strategy.
4.2.1 Key Observations
Across all tasks and models, we observe that, except for Temperature sampling, the other stochastic decoding methods achieve some of the highest SR ratios, maintaining good quality with minimal energy fluctuations across different hyperparameter choices. This suggests that these methods are less sensitive to hyperparameter tuning and are more energy-conscious relative to the quality they produce.
@@ 4.2.1 Key Observations
In most cases, especially in translation and coding tasks, beam-based strategies (BS and DBS) exhibit significant energy fluctuations despite minimal variation in quality. Conversely, in translation tasks in Qwen2.5 variants, Temperature sampling exhibits the opposite trend, with substantial quality variation but stable energy fluctuations. This shows that variations in quality and energy across hyperparameters do not always correlate.
@@
Another notable finding is that, while we identified AD as the most energy-saving method in our experiments, we observe that only in translation and coding tasks it ranks among the best methods with the highest SR ratio. In other cases, it exhibits variations in both generation quality and energy consumption across its selected hyperparameters, implying that AD performs optimally only when properly tuned in text generation tasks.
@@
Contrastive Search stands out as the most sensitive to hyperparameter changes among all methods, with $k=25$ exhibiting significantly great energy variations across different penalty $\alpha$ values. Larger $k$ values can lead to substantial shifts in both performance and energy consumption.
@@
Additionally, we observe a unique pattern in the DoLA method. While Table 2 shows that it is among the methods with the highest energy consumption, it also achieves some of the highest SR ratios in the Llama3.1-8B model. This indicates that even an energy-hungry decoding strategy may have minimal energy variation across different hyperparameters while still delivering acceptable (though not outstanding) generation quality in certain tasks, leading to high SR ratios.
Lastly, looking closely at the specific energy values for each hyperparameter, we observe that changing hyperparameters incrementally does not necessarily lead to a constant increase or decrease in energy. For example, when examining the average energy consumption of the min-p strategy across different hyperparameters in the HumanEval dataset with Qwen2.5-7B, we found that as the hyperparameter increases from 0.02 to 1.0, the energy fluctuates between 3.7 and 4.6 watt-hours, peaking at 5.4 watt-hours when the value is 0.05. This highlights that energy consumption does not always follow a smooth trend with hyperparameter changes, and small adjustments can sometimes lead to unexpected peaks or drops.
5 Energy–Latency Relationship
Since energy is calculated as the product of power and time, it depends not only on decoding latency but also on GPU power draw during measurement. When comparing two strategies, one with slightly higher runtime but lower power usage can still consume less energy overall, and vice versa. To check whether there is a difference between energy usage and wall-clock latency, we compared the energy ranking and latency ranking of decoding strategies across our tasks and models. We report the average decoding latency, average GPU power draw, and average GPU utilization for each decoding strategy in the Qwen2.5-7B-Instruct model in Table 11. The corresponding results for the Llama3.1-8B-Instruct and Qwen2.5-14B-Instruct models are reported in Tables 12 and 13 in Appendix.
As expected, models of different sizes or architectures draw different ranges of power for each decoding strategy. For example, in Qwen2.5-7B the average power draw for Greedy ranges between 109–111 watts across all tasks, while in Llama3.1-8B it ranges between 105–111 watts, and in Qwen2.5-14B it ranges between 120–125 watts. When looking at each task separately, we also observe different power draws for different decoding strategies, despite running the experiments at the same application clock frequency. This is related to differences in GPU utilization and the varying computational complexity of each strategy in each task. In general, stochastic methods show stable and high GPU utilization, leading to relatively similar power draw across methods. In contrast, deterministic strategies show large deviations in GPU utilization and power usage. For instance, in Qwen2.5-7B in translation task, Beam Search has 91% GPU utilization with a power draw of 102 watts, while DoLa reaches 96% utilization with 125 watts, and CS has 88% utilization with 103 watts of power draw.
To measure the relationship between energy rankings and latency rankings of different decoding strategies across various models and tasks, we use Kendall’s Tau correlation test. This non-parametric test is suitable because we have a relatively small sample size, the rankings do not follow a normal distribution, and the values are ordinal. The test shows how consistently energy and latency rankings are ordered in the same way across different cases. To better illustrate these results, we present them both as a heatmap (Fig 2), which highlights the correlation patterns, and as a table of p-values (Table 16 in Appendix).
Figure 2: Heatmap of Kendall’s Tau correlations between latency and energy consumption rankings of decoding strategies across different models and tasks.
Although in some cases, such as GSM8K and HumanEval with the Qwen2.5-7B and Llama3.1-8B models, the rankings are significantly correlated with minimal differences, in other cases—especially with Qwen2.5-14B—we observe smaller scores. For instance, in the German-to-English translation task, the test result is 0.64, compared to significantly higher values in other models. To better interpret these results, we analyze the energy and latency rankings. Interestingly, we find that Beam Search is one of the strategies with the largest deviation between its energy ranking and latency ranking. This deviation comes from its lower power draw of around 110 watts, compared to stochastic strategies that fluctuate around 120 watts. In this case, although Beam Search has higher latency, its lower power draw leads to lower overall energy consumption compared to strategies with lower latency but higher power draw. We observe similar situations in other cases where the energy rankings differ from the latency rankings of decoding strategies due to fluctuations in power draw. These observations show that energy consumption does not always align with latency due to differences in power draw among strategies. Understanding both is important for evaluating the overall efficiency of decoding methods.
@@ 6 Limitations & Future Work
In this study, we confined our experiments to Hugging Face/PyTorch rather than inference frameworks like vLLM[58] or SGLang[59]. When conducting our study, we found that these frameworks did not support some of the decoding strategies we wanted to evaluate. Thus, to focus on a larger subset of decoding strategies, we decided to leave testing these optimized inference frameworks for future studies. While we carefully tried to select a large range of hyperparameters for our experiments, our approach was not exhaustive, leaving a broader range of parameter settings yet to be explored. We also did not combine different decoding strategies, suggesting another avenue for future research. Finally, our analysis focused on GPU energy consumption; however, future work could extend this scope to encompass other factors such as CPU power usage, memory and temperature, providing a more comprehensive understanding of how decoding strategies influence overall resource consumption.
7 Conclusion
In this study, we investigated the impact of decoding strategies on both generative quality and GPU energy consumption during LLM inference across various text generation tasks. Through a comprehensive evaluation, we found that the choice of decoding strategy significantly affects GPU energy consumption during inference, even in cases where it has minimal impact on output quality. Additionally, we observed that different decoding methods present distinct trade-offs between text quality and energy efficiency. While deterministic methods such as Beam Search and DBS generally achieve top-tier quality in most generation tasks, they also tend to consume more energy and exhibit higher energy fluctuations across their hyperparameters. In contrast, stochastic methods appear to be more energy-efficient, often providing a better quality-to-energy ratio and demonstrating lower sensitivity to hyperparameter changes relative to their output quality. Overall, no single decoding strategy outperforms all others across every metric. Instead, the optimal choice depends on the specific task and whether the priority is maximizing generation quality, minimizing energy consumption, or achieving a stable balance between both.
From a practical standpoint, our findings suggest the following general recommendations: Assisted Decoding (AD) consistently achieves the lowest energy consumption across all tasks, making it suitable for energy-constrained applications. Stochastic sampling strategies such as Top-p and Min-p offer a robust trade-off between quality and energy efficiency, providing stable and balanced performance across different workloads. Conversely, beam-based deterministic methods like Beam Search and DBS deliver the highest overall quality but at the cost of significantly greater energy usage. Therefore, AI practitioners aiming for energy-efficient inference should prioritize AD or stochastic sampling, while those emphasizing generative accuracy may prefer beam-based approaches.
@@ 8 Ethical Considerations and Societal Impact
Code Availability
All the code used in this study is publicly available in our GitHub repository https://github.com/ds-anik/Energy-Conscious-LLM-Decoding.
Acknowledgments
The research presented in this paper has benefited from the Experimental Infrastructure for Exploration of Exascale Computing (eX3), which is financially supported by the Research Council of Norway under contract 270053.
@@ Appendix A Appendix
Table 4: Optimal Hyperparameters for the Decoding Strategies used in the experiments with Qwen2.5-7B-Instruct. The Greedy decoding method does not require any hyperparameters, so we represented it with a dash (-) in the table.